George Caloz & Frères, located in Grenchen, Switzerland, makes luxury custom watches in small lots. One of the company’s products, a platinum diving watch, goes through an etching process. The company has recorded etching costs as follows over the last six weeks:

Week Units Etching Cost

1 10 $ 27

2 7 $20

3 12 $31

4 6 $20

5 8 $24

6 11 $30

Total 54 $ 152

For planning purposes, management would like to know the variable etching cost per unit and the total fixed etching cost per week.

2-a. Using the least-squares regression method, estimate the variable etching cost per unit and the total fixed etching cost per week.

Variable etching Cost ____

Fixed Etching cost _____

2-b. Express these estimates in the form Y = a + bX. _______
Answer #1

2a.

x y xy x^2
10 27 270 100
7 20 140 49
12 31 372 144
6 20 120 36
8 24 192 64
11 30 330 121
54 152 1424 514
Unit Variable Cost (b) = nΣxy − (Σx)(Σy)
nΣx2 − (Σx)2
Variable cost = (6*1424) - (54*152) =   2
(6*514) - (54^2)
Total Fixed Cost (a) = Σy − bΣx / n
152 -( 2*54) / 6 = 7.33
2b.
Y = 7.33+ 2x